UNIVERSITY iianane 23426 MNa rr domain X-30-width 150.4 1400 130 1200 100 100. (0.0 50.0 40.0 』2n_0 10.0 @w Decouplsag x i parts per Millios 1 130


Assign all peaks and interpret spectra
Answer #1

10 он 2-(2-methoxyphenoxy)ethanol hydrate (Guaifenesin) 55 PPM = C-10, 63 PPM= C-1 70 PPM = C-3 71 PPM- C- 2, 1 1 1 PPM = C-6